# Twenty years of Nešetřil’s classification programme of Ramsey classes

### v3

@@ A.4 Recursive partite construction
By Observations A.13 and A.14, the usual construction used to prove the Picture lemma will yield $U$-closed embeddings only if $\mathbf{P}|_{\mathbf{A}}$ is $U$-closed, and this is typically true only if $\mathbf{A}$ is $U$-closed in $\mathbf{D}$. This is the main difficulty we will need to deal with, and in the following we need to carefully track which embeddings are guaranteed to be $U$-closed.
@@
Let $L$ be a language, let $\mathbf{A}$ and $\mathbf{D}$ be $L$-structures, let $\mathbf{B}$ be a $U$-transversal $\mathbf{D}$-partite system, and let $\alpha\colon\mathbf{A}\to\mathbf{D}$ be an embedding. Then there exists a $U$-transversal $\mathbf{D}$-partite system $\mathbf{C}$ such that for every $2$-colouring $\chi$ of $\mathop{\mathrm{Emb}}\nolimits(\mathbf{A},\mathbf{C})_{U,\alpha}$ there exists an embedding $f\colon\mathbf{B}\to\mathbf{C}$ such all embeddings $f\circ e$, $e\in\mathop{\mathrm{Emb}}\nolimits(\mathbf{A},\mathbf{B})_{U}$ are $U$-closed and have the same colour.
@@
Do the same construction as always, but use Lemma A.12 and then freely amalgamate copies of $\mathbf{B}$ only over $U$-closed embeddings $\mathbf{B}|_{\mathbf{A}}\to(\mathbf{B}|_{\mathbf{A}})^{N}$. Notice that Lemma A.12 promises us that $f$ can be chosen such that all embedding $f\circ e$ are $U$-closed in $(\mathbf{B}|_{\mathbf{A}})^{N}$. By Observation A.15 the process of extending $U$-closed embedding of $\mathbf{B}|_{\mathbf{A}}\to(\mathbf{B}|_{\mathbf{A}})^{N}$ to $\mathbf{B}$ by free amalgamation preserves this property, and so all embeddings $f\circ e$, $e\in\mathop{\mathrm{Emb}}\nolimits(\mathbf{A},\mathbf{B})_{U}$ are $U$-closed in $\mathbf{C}$.
The “moreover” part follows from Observations A.13 and A.14. ∎
@@
We will proceed analogously to the proofs of Theorems 2.6 and A.6. Fix $L$, $\mathbf{A}$, $\mathbf{B}$ as in the statement. Apply Theorem 2.5 to obtain $\mathbf{D}$ satisfying $\mathbf{D}\longrightarrow(\mathbf{B})^{\mathbf{A}}_{2}$. Let $\alpha_{0}$, $\alpha_{1}$, …, $\alpha_{m-1}$ be an enumeration of all (not necessarily $U$-closed) embeddings $\mathbf{A}\to\mathbf{D}$. We shall define $U$-transversal $\mathbf{D}$-partite systems $\mathbf{P}_{0},\mathbf{P}_{1},\ldots,\mathbf{P}_{m}$ such that for every $0\leq i<m$ and every 2-colouring $\chi$ of $\mathop{\mathrm{Emb}}\nolimits(\mathbf{A},\mathbf{P}_{i+1})_{U,\alpha_{i}}$ there is a $U$-closed embedding $f_{i}\colon\mathbf{P}_{i}\to\mathbf{P}_{i+1}$ such that $\chi$ is constant on $\{f_{i}\circ e:\mathop{\mathrm{Emb}}\nolimits(\mathbf{A},\mathbf{P}_{i})_{U,\alpha_{i}}\}$.

### v5

@@ A.4 Recursive partite construction
By Observations A.19 and A.20, the usual construction used to prove the Picture lemma will yield $U$-closed embeddings only if $\mathbf{P}|_{\mathbf{A}}$ is $U$-closed, and this is typically true only if $\mathbf{A}$ is $U$-closed in $\mathbf{D}$. This is the main difficulty we will need to deal with, and in the following we need to carefully track which embeddings are guaranteed to be $U$-closed.
@@
Let $L$ be a language, let $\mathbf{A}$ and $\mathbf{D}$ be $L$-structures, let $\mathbf{B}$ be a $U$-transversal $\mathbf{D}$-partite system, and let $\alpha\colon\mathbf{A}\to\mathbf{D}$ be an embedding. Then there exists a $U$-transversal $\mathbf{D}$-partite system $\mathbf{C}$ such that for every $2$-colouring $\chi$ of $\mathop{\mathrm{Emb}}\nolimits(\mathbf{A},\mathbf{C})_{U,\alpha}$ there exists an embedding $f\colon\mathbf{B}\to\mathbf{C}$ such that all embeddings $f\circ e$, $e\in\mathop{\mathrm{Emb}}\nolimits(\mathbf{A},\mathbf{B})_{U}$ are $U$-closed and have the same colour.
@@
Do the same construction as always, but use Lemma A.18 and then freely amalgamate copies of $\mathbf{B}$ only over $U$-closed embeddings $\mathbf{B}|_{\mathbf{A}}\to(\mathbf{B}|_{\mathbf{A}})^{N}$. Notice that Lemma A.18 promises us that $f$ can be chosen such that all embedding $f\circ e$ are $U$-closed in $(\mathbf{B}|_{\mathbf{A}})^{N}$. By Observation A.21 the process of extending $U$-closed embedding of $\mathbf{B}|_{\mathbf{A}}\to(\mathbf{B}|_{\mathbf{A}})^{N}$ to $\mathbf{B}$ by free amalgamation preserves this property, and so all embeddings $f\circ e$, $e\in\mathop{\mathrm{Emb}}\nolimits(\mathbf{A},\mathbf{B})_{U}$ are $U$-closed in $\mathbf{C}$.
The “moreover” part follows from Observations A.19 and A.20. ∎
@@
We will proceed analogously to the proofs of Theorems 2.9 and A.10. Fix $L$, $\mathbf{A}$, $\mathbf{B}$ as in the statement. Apply Theorem 2.7 to obtain $\mathbf{D}$ satisfying $\mathbf{D}\longrightarrow(\mathbf{B})^{\mathbf{A}}_{2}$. Let $\alpha_{0}$, $\alpha_{1}$, …, $\alpha_{m-1}$ be an enumeration of all (not necessarily $U$-closed) embeddings $\mathbf{A}\to\mathbf{D}$. We shall define $U$-transversal $\mathbf{D}$-partite systems $\mathbf{P}_{0},\mathbf{P}_{1},\ldots,\mathbf{P}_{m}$ such that for every $0\leq i<m$ and every 2-colouring $\chi$ of $\mathop{\mathrm{Emb}}\nolimits(\mathbf{A},\mathbf{P}_{i+1})_{U,\alpha_{i}}$ there is a $U$-closed embedding $f_{i}\colon\mathbf{P}_{i}\to\mathbf{P}_{i+1}$ such that $\chi$ is constant on $\{f_{i}\circ e:\mathop{\mathrm{Emb}}\nolimits(\mathbf{A},\mathbf{P}_{i})_{U,\alpha_{i}}\}$.